OpenStudy (anonymous):

How do i simplify -15arctan(3/4) + 15pi -20 to 15arctan(4/3) -20?

OpenStudy (anonymous):

\[-15\arctan(3/4)+15\pi-20 = 15(\pi-\arctan(3/4)-20\] \[=15(\pi-37°) -20\] \[15\arctan(-3/4)-20\] You cannot simplify it to the second equation

OpenStudy (jdoe0001):

wonder if there's a context or something

OpenStudy (anonymous):

Well basically, I have a function, which is 20x - 5x^2 / x^2 + 9 I have to find its primitive. Some part of it is the same as in the answers, but this part differs,so I'm trying to understand that...

OpenStudy (jdoe0001):

you may have better luck if maybe ask for the original material

OpenStudy (anonymous):

calculate the integral between 0 and 4 of the function i have provided above

OpenStudy (jdoe0001):

hmmm dunno it... try maybe posting anew for the integral instead, if someone knows, they'd help